### v1

# Perfect codes over non-prime power alphabets: an approach based on Diophantine equations

Abstract.
The classification of perfect codes over non-prime power alphabets has been an open problem for which there have been no new results in almost $50$ years. In this paper, we show non-existence of perfect $2-$error correcting codes over $q-$ary alphabets for more than 170 new values of $q$. Our methods rely on techniques from the resolution of generalised Ramanujan–Nagell equations and from computational number theory.
Key words and phrases:
Error correcting codes, $2-$perfect codes, Hamming bound, Ramanujan–Nagell equations, Mordell curves
2010 Mathematics Subject Classification:
Primary 94B65, 11D61, Secondary 11G05, 11G50, 14G05
1. Introduction
1.1. Background
Error correcting codes have been widely studied since the 1940s, both due to their mathematical interest and for their applicability in the transmission of information over noisy channels. In this paper, we shall consider the following standard definition of an error correcting code.
Definition 1.
Let $q\geq 2$, $e\geq 1$ be positive integers, and let $Z_{q}$ be a set with $q$ elements, which we shall call the alphabet. An $e-$error correcting code $C$ with parameters $(n,M)$ is a subset of $Z_{q}^{n}$ of size $M$ and such that any two elements $\overline{w}_{1},\overline{w}_{2}\in C$ differ in at least $12e+$ 1 positions.
It is a standard fact in coding theory (for example, see [10, Theorem 2.16]) that any $e-$error correcting code with parameters $(n,M)$ over an alphabet of size $q$ satisfies the sphere packing bound or Hamming bound:
$M\left(\sum_{j=0}^{e}{\binom{n}{j}}(q-1)^{j}\right)\leq q^{n}.$
In the case where the previous expression is an equality, we say that the code is perfect. In other words, $C$ is perfect if and only if $n$ and $M$ satisfy the following Diophantine equation:
(1) $M={q^{n}}\left/\sum_{j=0}^{e}{\binom{n}{j}}(q-1)^{j}\right..$
The classification of perfect codes was a very active topic of research in the late 20th century, and we refer the reader to [9] for an excellent survey on the most relevant results.
In particular, we highlight that if $q$ is a prime power, the parameters of $q-$ary perfect codes have been completely classified by Tietäväinen [22], and independently by Leontiev and Zinoniev [15].
On the other hand, if $q$ is not a prime power and $e\geq 3$, the work by Reuvers [19], Best [4] and Hong [11] shows that there are no perfect $e-$error correcting codes over any $q-$ary alphabet. In fact, many authors conjecture (see [9, Section 5], for example) the following:
Conjecture 1.
Let $q\geq 6$ be a non-prime power. Then, there are no perfect error correcting codes over $Z_{q}$.
While this question has been completely settled if $e\geq 3$, non-existence results are much more scarce if $e\leq 2$. For instance, if $e=2$, the work of Reuvers [19], van Lint [17] and Bassalygo et al. [2] show that perfect $2-$error correcting codes do not exist for
$q=6,10,15,21,22,26,30,35\quad\text{and}\quad q=2^{s}3^{r},\text{ where }s,r>0.$
However, these results date back to the 1970s, and it is surprising to note that there have been no new results since. In this paper, we address this gap in the literature by proving new non-existence results for the case $e=2$ and 172 new values of $q$, providing further evidence for Conjecture 1.
1.2. The main results
For the rest of the paper, we let $q\geq 6$ be a positive integer which is not a prime power. In this paper, we will prove new results related to the non-existence of perfect $2-$error correcting codes over alphabets of size $q$.
Since our methodology applies for $q=6,15,21,22,26,30,35$, for $q=10$, and for some cases where $q=2^{r}\cdot 3^{s}$, we shall recover the aforementioned results by Reuvers [19], van Lint [17] and Bassalygo et al [2].
Our main result is the following, which covers all but two values of $q\leq 200$, as well as some values of $q$ in the interval $q\in[201,600]$ subject to certain additional restrictions.
Theorem 2.
Let $q\geq 6$ be a positive integer which is not a prime power, satisfying any of the following conditions.
  1. (a)
    $q\leq 200$ and $q\neq 94,166$.
  2. (b)
    $q\leq 600$ and all prime divisors of $q$ are contained in the set $\{2,3,5,7,11\}$.
Then, there are no perfect $2-$error correcting codes over alphabets of size $q$.
The reason why we need to exclude $q=94$ and $q=166$ from Theorem 2 is due to computational difficulties. We shall comment more on this fact in Section 5.
The second result that we present applies to all values of $q$, but, instead of showing non-existence of perfect codes, it shows that there can only be finitely many perfect $2-$error correcting perfect codes over a $q-$ary alphabet. This is consistent with Conjecture 1, albeit much weaker.
Theorem 3.
Let $q\geq 6$ be a positive integer which is not a prime power. Then, there are at most finitely many perfect $2-$error correcting codes over an alphabet of size $q$.
In order to prove Theorems 2 and 3, we will study the Hamming bound (1) with $e=2$. We shall see that, for a fixed value of $q$, it can be reduced to a Ramanujan–Nagell type equation, which we can solve effectively with the use of several techniques from computational number theory. If, after resolving the equation, there are any solutions $(n,M)$ to (1), we can use Lloyd’s theorem to prove that they cannot constitute a perfect code.
The organisation of this paper is as follows. In Section 2, we will reduce the Hamming bound to a Ramanujan–Nagell equation and prove Theorem 3. In Section 3, we will build upon the work of von Känel and Matschke [12] in order to obtain all solutions to the previously obtained Ramanujan–Nagell equations. In Section 4, we will show that the outstanding solutions cannot form perfect codes by using Lloyd’s theorem, and therefore prove Theorem 2. Finally, in Section 5, we shall discuss complications arising from our methodology, as well as potential future lines of work.
All computations in this paper have been implemented in Magma [5] and are available for the reader in the author’s GitHub repository: https://github.com/PJCazorla/perfect-q-ary-codes.
Acknowledgements: The author would like to thank Yuri Bazlov for insightful discussions.
2. The Hamming bound and Ramanujan–Nagell type equations
We suppose that there exists a perfect 2 error-correcting $q-$ary code with parameters $(n,M)$. In order to avoid trivial codes, we shall assume that $n\geq 5$ and $M\geq 2$. Then, it follows from the Hamming bound (1) with $e=2$ that
(2) $\left(1+n(q-1)+\frac{n(n-1)}{2}(q-1)^{2}\right)M=q^{n}.$
Let $q_{1},\dots,q_{k}$ denote the prime divisors of $q$. By (2), $M\mid q^{n}$ and, consequently, there is a solution $(n,n_{1},\dots,n_{k})\in\mathbb{Z}^{k+1}$ to the Diophantine equation
(3) $1+n(q-1)+\frac{n(n-1)}{2}(q-1)^{2}=q_{1}^{n_{1}}\dots q_{k}^{n_{k}}.$
with $n\geq 5$ and where $q\geq 6$ is fixed. Now, let us introduce some notation. Let $x$ and $D_{q}$ be given by the following expressions:
(4) $x=2n(q-1)+3-q\quad\text{and}\quad D_{q}=8-(q-3)^{2}.$
Multiplying (3) by 8 and completing the square, we see that the existence of a perfect 2-error correcting code over a $q-$ary alphabet implies that there exists a solution $(x,n_{1},\dots,n_{k})\in\mathbb{Z}^{k+1}$ to the Diophantine equation
(5) $x^{2}+D_{q}=8q_{1}^{n_{1}}\dots q_{k}^{n_{k}}.$
Let $A$ denote the set of positive integers supported only on $\{q_{1},\dots,q_{k}\}$, so that
$A=\{q_{1}^{\alpha_{1}}\dots q_{k}^{\alpha_{k}}\mid\alpha_{1},\dots,\alpha_{k}\geq 0\}.$
Clearly, the Diophantine equation (5) can be rewritten as
(6) $x^{2}+b=cy,\quad x\in\mathbb{Z},\quad y\in A,$
where $b=D_{q}$ and $c=8$. This is an example of a generalised Ramanujan–Nagell equation, as considered by von Känel and Matschke in [12]. We can immediately use their work to prove Theorem 3.
Proof of Theorem 3.
By [12, Corollary K], it follows that there are at most finitely many solutions to any generalised Ramanujan–Nagell equation such as (6) and, consequently, there are at most finitely many solutions to (5), showing that there can only be finitely many perfect $2-$error correcting $q-$ary codes.
For the convenience of the reader, we shall give a different proof of the fact that (6) has finitely many solutions. This proof will allow us to fix notation and will be insightful for our work in Section 3.
Without loss of generality, we can assume that $\gcd(c,q_{1}\dots q_{k})=1$. This is because any powers of $q_{1},\dots,q_{k}$ present in $c$ can be absorbed into $y$ without altering the number of solutions of (6). Then, we can write
(7) $c=c_{0}c_{1}^{3}\quad\text{and}\quad y=y_{0}y_{1}^{3},$
where $c_{0},c_{1},y_{0},y_{1}\in\mathbb{Z}$ with $c_{0}$ and $y_{0}$ cubefree. We note here that
(8) $c_{0}y_{0}\in\{c_{0}q_{1}^{a_{1}}\dots q_{k}^{a_{k}}\mid 0\leq a_{1},\dots,a_{k}\leq 2\}.$
We define $d$ and $z$ by
(9) $d=c_{0}y_{0}\quad\text{ and }\quad z=c_{1}y_{1},$
so that (6) can be rewritten as
$x^{2}+b=dz^{3}.$
Let $(x,z)\in\mathbb{Z}^{2}$ be a solution to the previous equation and let $(X,Y)$ be given by
(10) $X=dz\quad\text{and}\quad Y=dx.$
Then, $(X,Y)$ is an integral point on the Mordell curve
(11) $E_{d}:Y^{2}=X^{3}-d^{2}b.$
By the work of Siegel [20], these curves have finitely many integral points. In addition, by (8), there are only finitely many curves $E_{d}$ to consider. Consequently, there are at most finitely many solutions to (6), finishing the proof. ∎
From the definition of $x$ and $D_{q}$ given in (4), we note that $x$ and $D_{q}$ are even precisely when $q$ is odd. Therefore, we can define
$x^{\prime}=\frac{x}{2}\quad\text{and}\quad D^{\prime}_{q}=\frac{D_{q}}{2}.$
In these cases, we can divide (5) by $4$ and obtain the following equation, which, computationally, is slightly easier to solve:
$(x^{\prime})^{2}+(D^{\prime}_{q})=2q_{1}^{n_{1}}\dots q_{k}^{n_{k}}.$
If we let $b=D^{\prime}_{q}$ and $c=2$, this is another instance of a generalised Ramanujan–Nagell equation in the form of (6). We shall present an algorithm for resolving these equations in Section 3.
3. Resolving the Diophantine equations
As we explained in Section 2, the existence of a perfect $2-$error correcting code over a $q-$ary alphabet implies the existence of a solution $(x,y)$ to the generalised Ramanujan–Nagell equation (6). In this section, we shall present an algorithm to solve this type of equations. Our algorithm is implemented in Magma in the GitHub repository https://github.com/PJCazorla/perfect-q-ary-codes and is based on [12, Algorithm 6.2].
As we discussed in the proof of Theorem 3, we can reduce the resolution of (6) to the determination of all integral points $(X,Y)$ on a finite number of Mordell curves $E_{d}$. Then, by (7), (9) and (10), we may recover the original solutions $x$ and $y$ via the expressions
$x=\frac{Y}{d}\quad\text{and}\quad y=\frac{d}{c_{0}}\left(\frac{X}{c_{1}d}\right)^{3}.$
In order to compute the integral points on the Mordell curves $E_{d}$, we distinguish two cases. Following [12, Section 4.1], we define $a_{S}$ by
$a_{S}=1728\prod_{p}p^{\min\{\text{ord}_{p}(bd^{2}),2\}},$
where the product is taken over all primes $p$ and $\text{ord}_{p}(.)$ denotes the standard $p-$adic valuation. If $a_{S}<500,000$, we can use [12, Algorithm 4.2] to compute all integral points very efficiently. This restriction is due to the fact that the algorithm requires the computation of all elliptic curves of conductor $N\mid a_{S}$. If $a_{S}<500,000$, these curves have been computed by Cremona [6] and so the necessary computational effort is minimal.
In the case where $a_{S}\geq 500,000$, our code computes the integral points by using Magma’s function IntegralPoints, which is based upon the use of linear forms in complex and $p$-adic elliptic logarithms (for a reference, see [21]). In this case, the main computational difficulty lies in the fact that a Mordell-Weil basis for $E_{d}$ needs to be computed which, in general, is very hard. We introduce two tricks to mitigate this difficulty.
Firstly, we note that the question of determining the rank $r$ of $E_{d}$ is highly non-trivial. However, by the work of Kolyvagin [13], we know that if the analytic rank of an elliptic curve is 00 or $1$, its algebraic rank coincides with its analytic rank. Since computing the analytic rank is much easier computationally than computing the algebraic rank, there are substantial computational savings in these cases.
We also note that if $r=0$, there is no need to compute any generators to the Mordell-Weil group. Similarly, if $r=1$, a generator can be found by computing Heegner points, as shown by Gross and Zagier [8]. We implement these two improvements in our code. While it could seem that these are not significant for the computation, they actually help tremendously and allow us to solve many equations which would otherwise not be amenable to our techniques.
Alternatively, we could have used the work of Bennett and Ghadermarzi [3], who computed all integral points on the Mordell curves
$E_{k}:y^{2}=x^{3}+k,$
where $0<|k|\leq 10^{7}$. They use a different set of techniques, involving the resolution of cubic Thue equations. This entrails a similar level of computational complexity to our approach, and it does not introduce any substantial computational improvements.
With the presented algorithm, we are able to solve the generalised Ramanujan–Nagell equations and recover the values of $n$ and $M$ which could represent a perfect $2-$error correcting $q-$ary code. This is the content of the following Lemma.
Lemma 3.1.
Let $q\geq 6$ be a positive integer which is not a prime power satisfying either of the two conditions in Theorem 2. Then, all solutions $(n,M)$ to (2) with $n\geq 5$ are given in Table 1.
Proof.
Using the algorithm that we have previously presented, we obtain all solutions $(x,y)\in\mathbb{Z}^{2}$ to each of the generalised Ramanujan–Nagell equations. By (4), we can then recover $n$ as
$n=\frac{x+q-3}{2(q-1)},$
and obtaining $M$ is immediate from (2). All the solutions that we found are contained in Table 1. ∎
$q$ $n$ $M$
$15$ $11$ $3^{4}\cdot 5^{10}$
$21$ $52$ $3^{40}\cdot 7^{52}$
$46$ $93$ $2^{79}\cdot 23^{91}$
Table 1. Solutions to (2) with $n\geq 5$ and $q$ satisfying the conditions in Theorem 2.
4. Dealing with the outstanding solutions: Lloyd’s theorem
In order to finish the proof of Theorem 2, we need to show that none of the solutions in Table 1 can be parameters for a perfect $q-$ary code. For this purpose, we shall use Lloyd’s Theorem, which was originally proved by Lloyd [18] if $q$ is a prime power. The general case was proved independently by Bassalygo [1], Delsarte [7] and Lenstra [14]. The statement of that theorem is as follows.
Theorem 4.
(Lloyd) Let $n,e$ and $q$ be integers. Suppose that there exists a perfect $e-$error correcting code over a $q-$ary alphabet with word length $n$. Then, the polynomial
$L_{e}(x)=\sum_{i=0}^{e}(-1)^{i}(q-1)^{e-i}\binom{x-1}{i}\binom{n-x}{e-i}$
has e distinct integer zeros in the interval $[1,n]$.
Our previous work, together with Lloyd’s theorem, allow us to finish the proof of Theorem 2.
Proof of Theorem 2.
Let $q$ be a number satisfying the conditions in the theorem and suppose for contradiction that there exists a perfect $2-$error correcting code over a $q-$ary alphabet, with parameters $(n,M)$.
By our discussion in Sections 2 and 3, along with Lemma 3.1, we know that the tuple $(q,n,M)$ is contained in one of the rows of Table 1. By Theorem 4, it follows that the polynomial $L_{2}(x)$ given by
$L_{2}(x)=(q-1)^{2}\frac{(n-x)(n-x-1)}{2}-(q-1)(x-1)(n-x)+\frac{(x-1)(x-2)}{2}$
has two distinct integer roots $x_{1},x_{2}$ with $1\leq x_{1}<x_{2}\leq n$. However, we check with Magma that this is not the case for any of the values of $q$ and $n$ in Table 1.
Consequently, for all the values of $q$ in the statement of the Theorem, there does not exist a perfect $2-$error correcting code over an alphabet of size $q$. ∎
5. Concluding remarks
We highlight that, with the methodology that we present here, it seems unlikely that the values of $q$ in Theorem 2 can be extended further.
This is due to the fact that, despite the computational improvements that we have introduced in Section 3, determining generators of the group $E_{d}(\mathbb{Q})$, where $E_{d}$ is the Mordell curve (11), is very hard computationally as the prime divisors of $q$ grow. For instance, while trying to solve the Ramanujan–Nagell equation arising from the case $q=94=2\cdot 47$, we need to find all integral points on the Mordell curve
$E_{2\cdot 47^{2}}:y^{2}=x^{3}+161478403652,$
which has conductor $N=16328471028588=2^{2}\cdot 3^{3}\cdot 47^{2}\cdot 8273^{2}$. We are able to show that $E_{2\cdot 47^{2}}$ has rank $1$, but the conductor is too large to apply the Heegner point methodology and, consequently, we are unable to find a generator of $E_{2\cdot 47^{2}}(\mathbb{Q})$.
More generally, our techniques will not be able to resolve the problem if $q$ has large prime divisors. For this reason, we have been able to solve many cases of $q>200$ by assuming that all prime divisors of $q$ lied in the set $\{2,3,5,7,11\}$. Even on this case, the necessary computational effort is significant, and so it is unlikely that this could be generalised to larger values of $q$.
Apart from the computational challenges, we remark that this methodology is not able to deal with families of $q$, but only with individual values. In order to prove Conjecture 1, even just for $e=2$, it seems likely that significantly new ideas will need to be introduced.
However, the methods that we present here could be used to approach perfect quantum codes, by mimicking this approach and changing the perfect code condition (1) by the perfect quantum code condition:
$M={q^{n}}\left/\sum_{j=0}^{e}{\binom{n}{j}}(q^{2}-1)^{j}\right..$
Li and Xing [16] classified perfect quantum codes over $q-$dimensional qudits, where $q$ is a prime power. In addition, they proved an analogue of Lloyd’s theorem for quantum codes. This makes the situation completely analogous to the classical case and, for this reason, it seems plausible that the approach that we presented here would directly generalise to quantum codes.